# MultiGraSCCo: A Multilingual Anonymization Benchmark with Annotations of Personal Identifiers

### 11 Mar 2026 v2

@@ Abstract
3Technical University Berlin, 4Humboldt University Berlin, 5University of Potsdam, 6BIFOLD
@@ 6.2.  Translation Quality
In the monolingual setting with mmBERT, we observe a strong baseline performance across all languages, illustrated in Table˜2. PHI detection is consistently high, with PHI micro F1-scores ranging from 0.90 to 0.95 and solid macro F1-scores across most languages. IPI detection, by contrast, remains more challenging. While all IPI micro F1-scores fall within a narrow range, the corresponding macro F1-scores are notably lower. This gap indicates that minority IPI classes are driving much of the overall error. These latter labels typically reflect more subtle, context-dependent cues, unlike PHI entities that often have explicit surface forms (e.g., names or dates). For completeness, we also report results from the language-specific models in the Appendix (Table˜8), which show similar trends but generally lower performance compared to mmBERT.
@@ 8.  Bibliographical References
  • I. Baroud, L. Raithel, S. Möller, and R. Roller (2025) Beyond de-identification: a structured approach for defining and detecting indirect identifiers in medical texts. In Proceedings of the Sixth Workshop on Privacy in Natural Language Processing, I. Habernal, S. Ghanavati, V. Jain, T. Igamberdiev, and S. Wilson (Eds.), Albuquerque, New Mexico, pp. 75–85. External Links: Link, Document, ISBN 979-8-89176-246-6 Cited by: §1, §2, §3.1, §3.
@@
  • R. Catelli, F. Gargiulo, V. Casola, G. De Pietro, H. Fujita, and M. Esposito (2020) Crosslingual named entity recognition for clinical de-identification applied to a covid-19 italian data set. Applied Soft Computing 97, pp. 106779. External Links: ISSN 1568-4946, Document, Link Cited by: §2.
@@
  • H. Dalianis, M. Duneld, and S. Velupillai (2009) The stockholm epr corpus–characteristics and some initial findings. Proceedings of 14th International Symposium for Health Information Management Research 219, pp. . Cited by: §2.
@@
  • H. Dalianis and S. Velupillai (2010) De-identifying swedish clinical text-refinement of a gold standard and experiments with conditional random fields. Journal of biomedical semantics 1 (1), pp. 6. External Links: Document Cited by: §2.
@@
  • A. Daza and A. Frank (2020) X-SRL: a parallel cross-lingual semantic role labeling dataset. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), B. Webber, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 3904–3914. External Links: Link, Document Cited by: §2.
@@
  • A. Feder, D. Vainstein, R. Rosenfeld, T. Hartman, A. Hassidim, and Y. Matias (2020) Active deep learning to detect demographic traits in free-form clinical notes. Vol. 107, pp. 103436. External Links: ISSN 1532-0464, Document, Link Cited by: §1, §2, §3.
@@
  • L. Hennig, P. Thomas, and S. Möller (2023) MultiTACRED: a multilingual version of the TAC relation extraction dataset. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 3785–3801. External Links: Link, Document Cited by: §2.
@@
  • G. Hripcsak and A. S. Rothschild (2005) Agreement, the f-measure, and reliability in information retrieval. Journal of the American medical informatics association 12 (3), pp. 296–298. Cited by: §$3.$
@@
  • A. E. Johnson, T. J. Pollard, L. Shen, L. H. Lehman, M. Feng, M. Ghassemi, B. Moody, P. Szolovits, L. Anthony Celi, and R. G. Mark (2016) MIMIC-iii, a freely accessible critical care database. Scientific data 3 (1), pp. 1–9. Cited by: §2.
@@
  • J. Klie, M. Bugert, B. Boullosa, R. E. de Castilho, and I. Gurevych (2018) The inception platform: machine-assisted and knowledge-oriented interactive annotation. In Proceedings of the 27th International Conference on Computational Linguistics: System Demonstrations, pp. 5–9 (en). Note: Event Title: The 27th International Conference on Computational Linguistics (COLING 2018) External Links: Link Cited by: §3.2.
  • V. Kocaman, Y. Mellah, H. Haq, and D. Talby (2023) Automated de-identification of Arabic medical records. In Proceedings of ArabicNLP 2023, H. Sawaf, S. El-Beltagy, W. Zaghouani, W. Magdy, A. Abdelali, N. Tomeh, I. Abu Farha, N. Habash, S. Khalifa, A. Keleg, H. Haddad, I. Zitouni, K. Mrini, and R. Almatham (Eds.), Singapore (Hybrid), pp. 33–40. External Links: Link, Document Cited by: §2, §4.2.
@@
  • T. Kocmi, R. Bawden, O. Bojar, A. Dvorkovich, C. Federmann, M. Fishel, T. Gowda, Y. Graham, R. Grundkiewicz, B. Haddow, R. Knowles, P. Koehn, C. Monz, M. Morishita, M. Nagata, T. Nakazawa, M. Novák, M. Popel, and M. Popovic (2022) Findings of the 2022 conference on machine translation (wmt22). In Conference on Machine Translation, External Links: Link Cited by: §4.3.
@@
  • T. Kolditz, C. Lohr, J. Hellrich, L. Modersohn, B. Betz, M. Kiehntopf, and U. Hahn (2019) Annotating german clinical documents for de-identification. Vol. 264, pp. 203–207. External Links: Document Cited by: §1, §2, §3.
@@
  • P. Lewis, B. Oguz, R. Rinott, S. Riedel, and H. Schwenk (2020) MLQA: evaluating cross-lingual extractive question answering. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, D. Jurafsky, J. Chai, N. Schluter, and J. Tetreault (Eds.), Online, pp. 7315–7330. External Links: Link, Document Cited by: §2.
@@
  • B. Li and C. Callison-Burch (2023) PAXQA: generating cross-lingual question answering examples at training scale. In Findings of the Association for Computational Linguistics: EMNLP 2023, H. Bouamor, J. Pino, and K. Bali (Eds.), Singapore, pp. 439–454. External Links: Link, Document Cited by: §2.
@@
  • J. Liu, Y. Lin, Z. Liu, and M. Sun (2019) XQA: a cross-lingual open-domain question answering dataset. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, A. Korhonen, D. Traum, and L. Màrquez (Eds.), Florence, Italy, pp. 2358–2368. External Links: Link, Document Cited by: §2.
@@
  • C. Lohr, F. Matthies, J. Faller, L. Modersohn, A. Riedel, U. Hahn, R. Kiser, M. Boeker, and F. Meineke (2024) De-identifying grascco–a pilot study for the de-identification of the german medical text project (gemtex) corpus. In German Medical Data Sciences 2024, pp. 171–179. Cited by: §1, §2, §2.
@@
  • M. Marimon, A. Gonzalez-Agirre, A. Intxaurrondo, H. Rodriguez, J. L. Martin, M. Villegas, and M. Krallinger (2019) Automatic de-identification of medical texts in spanish: the meddocan track, corpus, guidelines, methods and evaluation of results.. In IberLEF@ SEPLN, pp. 618–638. Cited by: §2.
@@
  • M. Marone, O. Weller, W. Fleshman, E. Yang, D. Lawrie, and B. V. Durme (2025) MmBERT: a modern multilingual encoder with annealed language learning. External Links: 2509.06888, Link Cited by: §5.1.
@@
  • L. Modersohn, S. Schulz, C. Lohr, and U. Hahn (2022) GRASCCO—the first publicly shareable, multiply-alienated german clinical text corpus. In German Medical Data Sciences 2022–Future Medicine: More Precise, More Integrative, More Sustainable!, pp. 66–72. Cited by: §1, §2, §2.
@@
  • M. Neves, A. Jimeno Yepes, A. Névéol, R. Bawden, G. M. Di Nunzio, R. Roller, P. Thomas, F. Vezzani, M. Vicente Navarro, L. Yeganova, D. Wiemann, and C. Grozea (2023) Findings of the WMT 2023 biomedical translation shared task: evaluation of ChatGPT 3.5 as a comparison system. In Proceedings of the Eighth Conference on Machine Translation, P. Koehn, B. Haddow, T. Kocmi, and C. Monz (Eds.), Singapore, pp. 43–54. External Links: Link, Document Cited by: §2.
@@
  • E. Nuutinen, I. Rastas, and F. Ginter (2025) Finnish SQuAD: A simple approach to machine translation of span annotations. In Proceedings of the Joint 25th Nordic Conference on Computational Linguistics and 11th Baltic Conference on Human Language Technologies (NoDaLiDa/Baltic-HLT 2025), R. Johansson and S. Stymne (Eds.), Tallinn, Estonia, pp. 424–432. External Links: Link, ISBN 978-9908-53-109-0 Cited by: §2.
@@
  • T. M. Seinen, J. A. Kors, E. M. van Mulligen, and P. R. Rijnbeek (2024) Annotation-preserving machine translation of english corpora to validate dutch clinical concept extraction tools. Journal of the American Medical Informatics Association 31 (8), pp. 1725–1734. Cited by: §2.
@@
  • A. Stubbs and Ö. Uzuner (2015) Annotating longitudinal clinical narratives for de-identification: the 2014 i2b2/uthealth corpus. Journal of Biomedical Informatics 58, pp. S20–S29. Note: Supplement: Proceedings of the 2014 i2b2/UTHealth Shared-Tasks and Workshop on Challenges in Natural Language Processing for Clinical Data External Links: ISSN 1532-0464, Document, Link Cited by: §2.
@@
  • L. Sweeney (2002) K-anonymity: a model for protecting privacy. International journal of uncertainty, fuzziness and knowledge-based systems 10 (05), pp. 557–570. Cited by: §1, §2.
@@
  • P. Tulajiang, Y. Sun, Y. Zhang, Y. Le, K. Xiao, and H. Lin (2025) A bilingual legal ner dataset and semantics-aware cross-lingual label transfer method for low-resource languages. ACM Trans. Asian Low Resour. Lang. Inf. Process. 24, pp. 103:1–103:21. External Links: Link Cited by: §2.
@@
  • H. Xu, A. Sharaf, Y. Chen, W. Tan, L. Shen, B. Van Durme, K. Murray, and Y. J. Kim (2024) Contrastive preference optimization: pushing the boundaries of llm performance in machine translation. In International Conference on Machine Learning, pp. 55204–55224. Cited by: §4.3.
@@
  • H. Zhan, D. C. Ong, and J. J. Li (2023) Evaluating subjective cognitive appraisals of emotions from large language models. In Findings of the Association for Computational Linguistics: EMNLP 2023, H. Bouamor, J. Pino, and K. Bali (Eds.), Singapore, pp. 14418–14446. External Links: Link, Document Cited by: §6.2.
@@ Appendix E Additional Training Details
  • German: learning rate 5e-5, batch size 8, epochs 8, weight decay $0.$
  • French, Persian, Italian, Polish, Russian, Turkish, and Ukrainian: learning rate 3e-5, batch size 8, epochs 6, weight decay $0.$
Language All PHI IPI
Micro Macro Micro Macro Micro Macro
English 0.859 0.756 0.933 0.878 0.774 0.572
French 0.819 0.539 0.907 0.729 0.710 0.253
German 0.717 0.516 0.822 0.621 0.606 0.358
Arabic 0.785 0.659 0.881 0.777 0.678 0.483
Persian 0.851 0.755 0.919 0.835 0.775 0.637
Italian 0.830 0.705 0.914 0.847 0.731 0.493
Polish 0.756 0.607 0.852 0.760 0.664 0.404
Russian 0.771 0.639 0.860 0.753 0.678 0.468
Ukrainian 0.631 0.504 0.784 0.674 0.480 0.250
Turkish 0.768 0.655 0.850 0.778 0.684 0.472
Table 8: Monolingual experiment results (language-specific models).

### 14 Mar 2026 v3

@@ Abstract
3Technische Universität Berlin, 4Humboldt University Berlin, 5University of Potsdam, 6BIFOLD
@@ 6.2.  Translation Quality
In the monolingual setting with mmBERT, we observe a strong baseline performance across all languages, illustrated in Table˜2. PHI detection is consistently high, with PHI micro F1-scores ranging from 0.90 to 0.95 and solid macro F1-scores across most languages. IPI detection, by contrast, remains more challenging. While all IPI micro F1-scores fall within a narrow range, the corresponding macro F1-scores are notably lower. This gap indicates that minority IPI classes are driving much of the overall error. These latter labels typically reflect more subtle, context-dependent cues, unlike PHI entities that often have explicit surface forms (e.g., names or dates).
@@ 8.  Bibliographical References
  • Baroud et al. (2025) Ibrahim Baroud, Lisa Raithel, Sebastian Möller, and Roland Roller. 2025. Beyond de-identification: A structured approach for defining and detecting indirect identifiers in medical texts. In Proceedings of the Sixth Workshop on Privacy in Natural Language Processing, pages 75–85, Albuquerque, New Mexico. Association for Computational Linguistics.
@@
  • Catelli et al. (2020) Rosario Catelli, Francesco Gargiulo, Valentina Casola, Giuseppe De Pietro, Hamido Fujita, and Massimo Esposito. 2020. Crosslingual named entity recognition for clinical de-identification applied to a covid-19 italian data set. Applied Soft Computing, 97:106779.
@@
  • Dalianis et al. (2009) Hercules Dalianis, Martin Duneld, and Sumithra Velupillai. 2009. The stockholm epr corpus–characteristics and some initial findings. Proceedings of 14th International Symposium for Health Information Management Research, 219.
@@
  • Dalianis and Velupillai (2010) Hercules Dalianis and Sumithra Velupillai. 2010. De-identifying swedish clinical text-refinement of a gold standard and experiments with conditional random fields. Journal of biomedical semantics, 1(1):6.
@@
  • Daza and Frank (2020) Angel Daza and Anette Frank. 2020. X-SRL: A parallel cross-lingual semantic role labeling dataset. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 3904–3914, Online. Association for Computational Linguistics.
@@
  • Feder et al. (2020) Amir Feder, Danny Vainstein, Roni Rosenfeld, Tzvika Hartman, Avinatan Hassidim, and Yossi Matias. 2020. Active deep learning to detect demographic traits in free-form clinical notes. volume 107, page 103436.
@@
  • Hennig et al. (2023) Leonhard Hennig, Philippe Thomas, and Sebastian Möller. 2023. MultiTACRED: A multilingual version of the TAC relation extraction dataset. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 3785–3801, Toronto, Canada. Association for Computational Linguistics.
@@
  • Hripcsak and Rothschild (2005) George Hripcsak and Adam S Rothschild. 2005. Agreement, the f-measure, and reliability in information retrieval. Journal of the American medical informatics association, 12(3):296–298.
@@
  • Johnson et al. (2016) Alistair EW Johnson, Tom J Pollard, Lu Shen, Li-wei H Lehman, Mengling Feng, Mohammad Ghassemi, Benjamin Moody, Peter Szolovits, Leo Anthony Celi, and Roger G Mark. 2016. Mimic-iii, a freely accessible critical care database. Scientific data, 3(1):1–9.
@@
  • Klie et al. (2018) Jan-Christoph Klie, Michael Bugert, Beto Boullosa, Richard Eckart de Castilho, and Iryna Gurevych. 2018. The inception platform: Machine-assisted and knowledge-oriented interactive annotation. In Proceedings of the 27th International Conference on Computational Linguistics: System Demonstrations, pages 5–9. Association for Computational Linguistics. Event Title: The 27th International Conference on Computational Linguistics (COLING 2018).
  • Kocaman et al. (2023) Veysel Kocaman, Youssef Mellah, Hasham Haq, and David Talby. 2023. Automated de-identification of Arabic medical records. In Proceedings of ArabicNLP 2023, pages 33–40, Singapore (Hybrid). Association for Computational Linguistics.
@@
  • Kocmi et al. (2022) Tom Kocmi, Rachel Bawden, Ondrej Bojar, Anton Dvorkovich, Christian Federmann, Mark Fishel, Thamme Gowda, Yvette Graham, Roman Grundkiewicz, Barry Haddow, Rebecca Knowles, Philipp Koehn, Christof Monz, Makoto Morishita, Masaaki Nagata, Toshiaki Nakazawa, Michal Novák, Martin Popel, and Maja Popovic. 2022. Findings of the 2022 conference on machine translation (wmt22). In Conference on Machine Translation.
@@
  • Kolditz et al. (2019) Tobias Kolditz, Christina Lohr, Johannes Hellrich, Luise Modersohn, Boris Betz, Michael Kiehntopf, and Udo Hahn. 2019. Annotating german clinical documents for de-identification. volume 264, pages 203–207.
@@
  • Lewis et al. (2020) Patrick Lewis, Barlas Oguz, Ruty Rinott, Sebastian Riedel, and Holger Schwenk. 2020. MLQA: Evaluating cross-lingual extractive question answering. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7315–7330, Online. Association for Computational Linguistics.
@@
  • Li and Callison-Burch (2023) Bryan Li and Chris Callison-Burch. 2023. PAXQA: Generating cross-lingual question answering examples at training scale. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 439–454, Singapore. Association for Computational Linguistics.
@@
  • Liu et al. (2019) Jiahua Liu, Yankai Lin, Zhiyuan Liu, and Maosong Sun. 2019. XQA: A cross-lingual open-domain question answering dataset. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 2358–2368, Florence, Italy. Association for Computational Linguistics.
@@
  • Lohr et al. (2024) Christina Lohr, Franz Matthies, Jakob Faller, Luise Modersohn, Andrea Riedel, Udo Hahn, Rebekka Kiser, Martin Boeker, and Frank Meineke. 2024. De-identifying grascco–a pilot study for the de-identification of the german medical text project (gemtex) corpus. In German Medical Data Sciences 2024, pages 171–179. IOS Press.
@@
  • Marimon et al. (2019) Montserrat Marimon, Aitor Gonzalez-Agirre, Ander Intxaurrondo, Heidy Rodriguez, Jose Lopez Martin, Marta Villegas, and Martin Krallinger. 2019. Automatic de-identification of medical texts in spanish: the meddocan track, corpus, guidelines, methods and evaluation of results. In IberLEF@ SEPLN, pages 618–638.
@@
  • Marone et al. (2025) Marc Marone, Orion Weller, William Fleshman, Eugene Yang, Dawn Lawrie, and Benjamin Van Durme. 2025. mmbert: A modern multilingual encoder with annealed language learning.
@@
  • Modersohn et al. (2022) Luise Modersohn, Stefan Schulz, Christina Lohr, and Udo Hahn. 2022. Grascco—the first publicly shareable, multiply-alienated german clinical text corpus. In German Medical Data Sciences 2022–Future Medicine: More Precise, More Integrative, More Sustainable!, pages 66–72. IOS Press.
@@
  • Neves et al. (2023) Mariana Neves, Antonio Jimeno Yepes, Aurélie Névéol, Rachel Bawden, Giorgio Maria Di Nunzio, Roland Roller, Philippe Thomas, Federica Vezzani, Maika Vicente Navarro, Lana Yeganova, Dina Wiemann, and Cristian Grozea. 2023. Findings of the WMT 2023 biomedical translation shared task: Evaluation of ChatGPT 3.5 as a comparison system. In Proceedings of the Eighth Conference on Machine Translation, pages 43–54, Singapore. Association for Computational Linguistics.
@@
  • Nuutinen et al. (2025) Emil Nuutinen, Iiro Rastas, and Filip Ginter. 2025. Finnish SQuAD: A simple approach to machine translation of span annotations. In Proceedings of the Joint 25th Nordic Conference on Computational Linguistics and 11th Baltic Conference on Human Language Technologies (NoDaLiDa/Baltic-HLT 2025), pages 424–432, Tallinn, Estonia. University of Tartu Library.
@@
  • Seinen et al. (2024) Tom M Seinen, Jan A Kors, Erik M van Mulligen, and Peter R Rijnbeek. 2024. Annotation-preserving machine translation of english corpora to validate dutch clinical concept extraction tools. Journal of the American Medical Informatics Association, 31(8):1725–1734.
@@
  • Stubbs and Uzuner (2015) Amber Stubbs and Özlem Uzuner. 2015. Annotating longitudinal clinical narratives for de-identification: The 2014 i2b2/uthealth corpus. Journal of Biomedical Informatics, 58:S20–S29. Supplement: Proceedings of the 2014 i2b2/UTHealth Shared-Tasks and Workshop on Challenges in Natural Language Processing for Clinical Data.
@@
  • Sweeney (2002) Latanya Sweeney. 2002. k-anonymity: A model for protecting privacy. International journal of uncertainty, fuzziness and knowledge-based systems, 10(05):557–570.
@@
  • Tulajiang et al. (2025) Paerhati Tulajiang, Yuanyuan Sun, Yuanyu Zhang, Yingying Le, Kelaiti Xiao, and Hongfei Lin. 2025. A bilingual legal ner dataset and semantics-aware cross-lingual label transfer method for low-resource languages. ACM Trans. Asian Low Resour. Lang. Inf. Process., 24:103:1–103:21.
@@
  • Xu et al. (2024) Haoran Xu, Amr Sharaf, Yunmo Chen, Weiting Tan, Lingfeng Shen, Benjamin Van Durme, Kenton Murray, and Young Jin Kim. 2024. Contrastive preference optimization: Pushing the boundaries of llm performance in machine translation. In International Conference on Machine Learning, pages 55204–55224. PMLR.
@@
  • Zhan et al. (2023) Hongli Zhan, Desmond C. Ong, and Junyi Jessy Li. 2023. Evaluating subjective cognitive appraisals of emotions from large language models. In Findings of the Association for Computational Linguistics: EMNLP 2023, pages 14418–14446, Singapore. Association for Computational Linguistics.
@@ Appendix E Additional Training Details
  • German: learning rate 5e-5, batch size 8, epochs 8, weight decay 0.0.
  • French, Persian, Italian, Polish, Russian, Turkish, and Ukrainian: learning rate 3e-5, batch size 8, epochs 6, weight decay 0.0.